OpenStudy (anonymous):

The function f(x)=ksin6x models a continuous probability for 0

OpenStudy (anonymous):

\[0 \le x \le \pi/6\]

OpenStudy (nincompoop):

hmmm how do solve this?

OpenStudy (anonymous):